A body of weight W = 120N is pushed by a horizontal for P = t2 + 6. A frictional force f= t3- 5 resists
its motion. If the body starts from rest, find the velocity of the body after 3 seconds.
1
Expert's answer
2020-06-10T07:51:26-0400

"F=ma=m\frac{dv}{dt}\to v=\frac{1}{m}\int F(t)dt"


"v=\frac{1}{m}\int (t^2+6-t^3+5)dt=\frac{1}{m}(t^3/3+6t-t^4/4+5t)+C="


"=\frac{1}{m}(t^3/3+11t-t^4/4)+C"


if "t=0" "v=0 \to" "C=0"


We have


"v=\frac{1}{m}(t^3/3+11t-t^4/4)=\frac{1}{120/9.81}(3^3/3+11\cdot 3-3^4/4)=1.78 m/s"